Two balls weighing 2.5 kg and 5 kg are subject to an elastic collision. Before the collision, the first and lighter ball was at rest, and the second one was moving at a speed of 3.5 m/s. Determine the kinetic energy of a lighter ball after a collision. (Answer: V1'=3.66 m/s, V2' = 1.166 m/s and Ek1'= 16.7 J)


1
Expert's answer
2021-04-19T07:22:00-0400

Let "v_1 = 0, v_2 = 3.5m\/s" be the speeds of the balls before the collision, "m_1 = 2.5kg, m_2 = 2m_1 = 5kg" the masses of the balls, "v_1', v_2'" the speeds of the balls after the collision.

Suppose the collision was head-on and ellastic. Thus, according to the momentum conservation law, have:


"m_1v_1 + m_2v_2 = m_1v_1' + m_2v_2'"

Since "v_1 = 0":


"m_2v_2 = m_1v_1' + m_2v_2'\\
2m_1v_2 = m_1v_1' + 2m_1v_2'\\
2v_2 = v_1' + 2v_2'"

According to the energy conservation law:


"\dfrac{m_2v_2^2}{2} = \dfrac{m_1v_1'^2}{2} + \dfrac{m_2v_2'^2}{2}\\
\dfrac{2m_1v_2^2}{2} = \dfrac{m_1v_1'^2}{2} + \dfrac{2m_1v_2'^2}{2}\\
2v_2^2 = v_1'^2 + 2v_2'^2"

Expressing "v_1'" from "2v_2 = v_1' + 2v_2'" and substituting it into "2v_2^2 = v_1'^2 + 2v_2'^2", obtain:


"3v_2'^2 - 4v_2v_2' + v_2^2 = 0"

Substituting the values and solving the quadratic equation, obtain:


"v_2' = 1.166 m/s\space \text{or}\space v_2' = 3.5 m/s"

The last values in impossible, since it would mean that the first ball remains at rest after the collsion. Thus, subsituting "v_2' = 1.166 m\/s" into "2v_2 = v_1' + 2v_2'", obtain:


"v_1' = 2(v_2 - v_2') = 4.66m\/s"

And the kinetic energy is:


"E_1' = \dfrac{m_1v_1'^2}{2} = 27.2J"

Answer. 27.2J